OpenStudy (anonymous):

Find the derivative of g(x) = (1/x) at x=3 algebraically.

OpenStudy (anonymous):

Algebraically? Do you mean using the limit definition?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

\[\begin{align*}\frac{d}{dx}\ln x&=\lim_{h\to0}\frac{\ln(x+h)-\ln x}{h}\\ &=\lim_{h\to0}\frac{\ln\left(\frac{x+h}{x}\right)}{h}\\ &=\lim_{h\to0}\frac{1}{h}\ln\left(1+\frac{h}{x}\right) \end{align*}\] Substitute \(t=\dfrac{h}{x}\). Then, as \(h\to0\), you have \(t\to0\). This gives you \(h=xt\), or \(\dfrac{1}{h}=\dfrac{1}{xt}\). \[\begin{align*}\frac{d}{dx}\ln x&=\lim_{t\to0}\frac{1}{xt}\ln\left(1+t\right)\\ &=\frac{1}{x}\color{red}{\lim_{t\to0}\ln(1+t)^{1/t}} \end{align*}\] The red part is the limit definition of \(\ln e\), or just 1. So you're left with \(\dfrac{1}{x}\), the derivative of \(\ln x\).